OpenStudy (anonymous):

Write an equation in slope intercept form y=mx+b with points (6,7) with a slope -2

OpenStudy (anonymous):

y = 4x -25

OpenStudy (freckles):

point-slope form is: \[y-y_1=m(x-x_1) \text{ you are given } m=-2 \text{ and } (x_1,y_1)=(6,7)\]

OpenStudy (freckles):

plug in

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (freckles):

then write it in y=mx+b form by doing a little multiplication and addition

OpenStudy (anonymous):

\[y1-y = m(x1-x)\] m = slope....y1 and x1 equal point on the line

OpenStudy (anonymous):

Thanks